# KEPLA: A Knowledge-Enhanced Deep Learning Framework for Accurate Protein-Ligand Binding Affinity Prediction

### v3

@@ Abstract
Accurate prediction of protein-ligand binding affinity is critical for drug discovery. While recent deep learning approaches have demonstrated promising results, they often rely solely on structural features of proteins and ligands, overlooking their valuable biochemical knowledge associated with binding affinity. To address this limitation, we propose KEPLA, a novel deep learning framework that explicitly integrates prior knowledge from Gene Ontology and ligand properties to enhance prediction performance. KEPLA takes protein sequences and ligand molecular graphs as input and optimizes two complementary objectives: (1) aligning global representations with knowledge graph relations to capture domain-specific biochemical insights, and (2) leveraging cross attention between local representations to construct fine-grained joint embeddings for prediction. Experiments on two benchmark datasets across both in-domain and cross-domain scenarios demonstrate that KEPLA consistently outperforms state-of-the-art baselines. Furthermore, interpretability analyses based on knowledge graph relations and cross attention maps provide valuable insights into the underlying predictive mechanisms.
Introduction
The prediction of protein-ligand binding affinity (PLA) has been widely considered as one of the most important tasks in computational drug discovery [1]. Here, ligands are usually drug candidates, including small molecules and biologics, which can interact with proteins as agonists or inhibitors in biological processes to cure diseases. Given a protein, we want to understand how well a ligand can interact with this protein. The strength of interaction between them can be quantified as a numerical score (called the binding affinity), which potentially determines whether a ligand can effectively influence the protein (for example, to inactivate a protein to cure a disease). Although experimental methods [2] can measure binding affinity, these biological tests are laborious and time-consuming. Thus, in silico predictive approaches have become increasingly necessary and received widespread attention [3]. With in silico approaches, binding affinities can be quickly predicted in the early stage of drug discovery. The predicted binding affinities can help rank drug candidates and prioritize the appropriate ones for more precise subsequent tests to accelerate the drug screening process, instead of applying costly biological methods directly to screen the entire pool of numerous candidate molecules [4].
@@ Introduction
Results
Overview of KEPLA
The proposed KEPLA framework is illustrated in Figure 1. Given an input protein-ligand pair, we first encode the protein amino acid sequence and the ligand molecular graph separately using ESM and GCN blocks, respectively. To incorporate relevant biochemical knowledge into the representation learning process, the encoded representations are further embedded into a KG module and optimized with the KG embedding objective. Simultaneously, a cross attention module is employed to capture fine-grained interactions between the protein and ligand representations. This module includes a local interaction mapping step followed by a cross attention computation step, which together generate a joint protein-ligand representation. Finally, a fully connected MLP decoder produces a predictive score representing the binding affinity. Details of each module in the KEPLA model, along with the training procedure, are provided in the Methods section.
@@ Results
Figure 1: Overview of the KEPLA framework. (a) Strategy of existing PLA prediction models. (b) Strategy of our KEPLA model. The key difference lies in the incorporation of a biochemical KG, which enhances both protein and ligand encoding. (c) Detailed architecture of KEPLA. The model employs two separate encoders: an ESM-based encoder processes the amino acid sequence, and a GCN-based encoder handles the SMILES string. These generate both local and global representations for proteins and ligands. The global representations are used in the KG embedding module, while the local representations are input into a cross attention module for PLA prediction. (d) The cross attention module. Here, an interaction map is computed via the multiplication of protein and ligand local representations, which serves as a similarity matrix to derive cross attention. The cross attention enables fine-grained fusion of protein and ligand local representations, resulting in a joint representation. This joint representation is passed through an MLP decoder to predict binding affinity. (e) The KG embedding module. Here, linear transformations project the global protein and ligand representations into a semantic space. Combined with entity and relation embeddings from the KG, a score function evaluates the plausibility of triplets, which forms the objective of this module.
Evaluation strategies and metrics
We evaluate prediction performance on two public datasets, PDBbind [23] and CSAR-HiQ [24], using two distinct data split strategies to simulate in-domain and cross-domain scenarios. For in-domain evaluation, following [25], we use the refined set of PDBbind as the primary training data, as there is substantial overlap between the full general set and CSAR-HiQ dataset. We randomly split the protein-ligand complexes in the refined set into training and validation sets at a 9:1 ratio. For testing, we use the PDBbind core set and the CSAR-HiQ dataset after removing any overlapping complexes from the refined set. For cross-domain evaluation, we propose a clustering-based pair split strategy to simulate domain shift. We conduct cross-domain evaluation on the larger-scale PDBbind datasets. Specifically, we first use the single-linkage algorithm to cluster proteins and ligands by PSC (pseudo-amino acid composition) [26] and ECFP4 (extended connectivity fingerprint, up to four bonds) [27], respectively. After that, we randomly select 60% protein clusters and 60% ligand clusters from the clustering result, and consider all protein-ligand pairs between the selected ligands and proteins as source domain data. All the pairs between proteins and ligands in the remaining clusters are considered to be target domain data. Detailed clustering implementation is provided in the Supplementary Information. Under this split, source and target domains are disjoint and follow different distributions. In line with standard domain adaptation settings, we use all labeled source domain data and 80% of the unlabeled target domain data for training, where this 80% also serves as the validation set. The remaining 20% of labeled target data is used for testing. Compared to in-domain random splits, this cross-domain setup poses a greater challenge but better reflects real-world generalization scenarios in drug discovery.
To comprehensively evaluate the model performance, following [28, 29], we adopt Root Mean Square Error (RMSE), Mean Absolute Error (MAE), Pearson’s correlation coefficient (R), and the standard deviation (SD) in regression to measure the prediction error. Detailed descriptions of these metrics are provided in the Supplementary Information. For evaluation, lower RMSE, MAE, and SD values indicate better performance, whereas higher R values are preferable.
In-domain performance comparison
We compare KEPLA against a range of representative interaction-free and interaction-based baseline models under the random split setting. The interaction-free baselines include RF-Score [25], DeepDTA [10], GraphDTA (including GCN, GAT, GIN, and GAT-GCN) [30], DrugBAN [31]. The interaction-based baselines include Pafnucy [28], OnionNet [29], SGCN [32], GraphTrans [33], NL-GCN [34], GNN-DTI [35], DMPNN [36], MAT [37], DimeNet [38], CMPNN [39], SIGN [40], and GIANT [41]. All baseline models are implemented using the official source codes and are evaluated on the same training, validation, and test sets. The validation set is used to select the model checkpoint with the lowest RMSE for final testing. It is worth noting that some baselines are not originally developed for PLA prediction and have been adapted accordingly for a fair comparison. The detailed description and setup of the baselines can be found in the Supplementary Information.
Table 1 presents the quantitative results. We report the mean and standard deviation of four evaluation metrics across five independent runs. When comparing model categories, we observe that interaction-based models generally outperform interaction-free models, highlighting the advantage of incorporating 3D structural information and physical interactions in PLA prediction. Despite being an interaction-free method, our proposed KEPLA achieves the best performance across all experimental settings on both datasets, surpassing even all interaction-based baselines. Specifically, KEPLA improves RMSE by 5.28% on the PDBbind dataset and by 12.42% on the CSAR-HiQ dataset compared to the best-performing baseline GIANT. These results strongly support KEPLA’s superior predictive capability. Considering KEPLA requires only the molecular graph of ligand compounds and primary sequences of target proteins as input like most interaction-free methods, the outstanding performance benefits from knowledge-enhanced representation learning and local interaction modeling.
Table 1: In-domain performance comparison on the PDBbind and CSAR-HiQ datasets under random split (statistics over five independent runs)
Method PDBbind core set
RMSE $\downarrow$ MAE $\downarrow$ SD $\downarrow$ $\uparrow$
Interaction-free methods RF-Score 1.446 (0.008) 1.161 (0.007) 1.335 (0.010) 0.789 (0.003)
DeepDTA 1.639 (0.026) 1.315 (0.023) 1.689 (0.025) 0.718 (0.014)
GCN 1.735 (0.034) 1.343 (0.037) 1.719 (0.027) 0.613 (0.016)
GAT 1.765 (0.026) 1.354 (0.033) 1.740 (0.027) 0.601 (0.016)
GIN 1.640 (0.044) 1.261 (0.044) 1.621 (0.036) 0.667 (0.018)
GAT-GCN 1.562 (0.022) 1.191 (0.016) 1.558 (0.018) 0.697 (0.008)
DrugBAN 1.378 (0.025) 1.102 (0.019) 1.328 (0.020) 0.788 (0.015)
Interaction-based methods Pafnucy 1.585 (0.013) 1.284 (0.021) 1.563 (0.022) 0.695 (0.011)
OnionNet 1.407 (0.034) 1.078 (0.028) 1.391 (0.038) 0.768 (0.014)
SGCN 1.583 (0.033) 1.250 (0.036) 1.582 (0.320) 0.686 (0.015)
GraphTrans 1.539 (0.044) 1.182 (0.046) 1.521 (0.042) 0.714 (0.019)
NL-GCN 1.516 (0.019) 1.198 (0.013) 1.511 (0.024) 0.720 (0.010)
GNN-DTI 1.492 (0.025) 1.192 (0.032) 1.471 (0.051) 0.736 (0.021)
DMPNN 1.493 (0.016) 1.188 (0.009) 1.489 (0.014) 0.729 (0.006)
MAT 1.457 (0.037) 1.154 (0.037) 1.445 (0.033) 0.747 (0.013)
DimeNet 1.453 (0.027) 1.138 (0.026) 1.434 (0.023) 0.752 (0.010)
CMPNN 1.408 (0.028) 1.117 (0.031) 1.399 (0.025) 0.765 (0.009)
SIGN 1.316 (0.031) 1.027 (0.025) 1.312 (0.035) 0.797 (0.012)
GIANT 1.269 (0.020) 0.999 (0.018) 1.265 (0.024) 0.814 (0.008)
Ours KEPLA 1.202 (0.017) 0.936 (0.013) 1.195 (0.016) 0.832 (0.005)
Method CSAR-HiQ set
RMSE $\downarrow$ MAE $\downarrow$ SD $\downarrow$ $\uparrow$
Interaction-free methods RF-Score 1.947 (0.012) 1.466 (0.009) 1.796 (0.020) 0.723 (0.007)
DeepDTA 1.983 (0.036) 1.510 (0.059) 2.037 (0.051) 0.633 (0.036)
GCN 2.324 (0.079) 1.732 (0.065) 2.302 (0.061) 0.464 (0.047)
GAT 2.213 (0.053) 1.651 (0.061) 2.215 (0.050) 0.524 (0.032)
GIN 2.158 (0.074) 1.624 (0.058) 2.156 (0.088) 0.558 (0.047)
GAT-GCN 1.980 (0.055) 1.493 (0.046) 1.969 (0.057) 0.653 (0.026)
DrugBAN 1.672 (0.034) 1.334 (0.042) 1.728 (0.059) 0.767 (0.039)
Interaction-based methods Pafnucy 1.939 (0.103) 1.562 (0.094) 1.885 (0.071) 0.686 (0.027)
OnionNet 1.927 (0.071) 1.471 (0.031) 1.877 (0.097) 0.690 (0.040)
SGCN 1.902 (0.063) 1.472 (0.067) 1.891 (0.077) 0.686 (0.030)
GraphTrans 1.950 (0.072) 1.508 (0.069) 1.886 (0.083) 0.687 (0.033)
NL-GCN 1.840 (0.024) 1.393 (0.016) 1.817 (0.028) 0.716 (0.011)
GNN-DTI 1.972 (0.061) 1.547 (0.058) 1.834 (0.090) 0.709 (0.035)
DMPNN 1.886 (0.026) 1.488 (0.054) 1.865 (0.035) 0.697 (0.013)
MAT 1.879 (0.065) 1.435 (0.058) 1.816 (0.083) 0.715 (0.030)
DimeNet 1.805 (0.036) 1.338 (0.026) 1.798 (0.027) 0.723 (0.010)
CMPNN 1.839 (0.096) 1.411 (0.064) 1.767 (0.103) 0.730 (0.052)
SIGN 1.735 (0.031) 1.327 (0.040) 1.709 (0.044) 0.754 (0.014)
GIANT 1.666 (0.024) 1.242 (0.030) 1.633 (0.034) 0.779 (0.011)
Ours KEPLA 1.459 (0.019) 1.132 (0.026) 1.517 (0.032) 0.813 (0.012)
Cross-domain performance comparison
In-domain prediction under a random split setting is relatively easier and holds limited practical significance. Therefore, we further investigate the more realistic and challenging task of cross-domain PLA prediction, where the training and test data follow different distributions. To simulate this scenario, we partition the original PDBbind dataset into source and target domains using the clustering-based pair split strategy. As a complementary evaluation, we also perform experiments using a cold pair split on the PDBbind dataset. Specifically, we randomly select 70% of the proteins and ligands and collect all associated protein-ligand pairs as the training set. The remaining 30% of proteins and ligands are used to form the validation and test sets in a 3:7 ratio. This cold pair split strategy ensures that none of the proteins or ligands in the test set have been seen during training, providing a stringent test of model generalization.
Ablation study
We conduct an ablation study to investigate the contributions of two key components in KEPLA: the KG and the cross attention modules. The results of various KEPLA variants are presented in Figure 3. To evaluate the impact of KG enhancement, we consider three variants of KEPLA with different configurations: KEPLA using only the protein-side KG, KEPLA using only the ligand-side KG, and KEPLA without KG enhancement. As shown in Figure 3a, the model’s strong performance is clearly supported by the incorporation of external knowledge, validating the effectiveness of KG enhancement. Notably, the protein-side KG contributes more significantly to affinity prediction than the ligand-side counterpart. A likely explanation is that the protein KG contains a greater number of entities and richer relational information, offering more comprehensive context for representation learning.
Structural-level interpretability with cross attention
[…]
For complex 3UI7 (pyrazoloquinoline bound to human PDE10A), the ligand molecule pyrazoloquinoline features a fused structure composed of a pyrazole ring and a quinoline ring. The pyrazole ring is a five-membered nitrogen-containing heterocycle, while the quinoline ring consists of a pyridine ring fused with a benzene ring. Our model correctly identifies the nitrogen-containing heterocyclic region as critical for protein-ligand binding: the nitrogen atom bearing a hydrogen in the pyrazole ring acts as a hydrogen bond donor to the side chain of Gln729; the other nitrogen atom interacts with Thr685 via a solvent molecule; and the nitrogen atom in the pyridine ring serves as a hydrogen bond acceptor to Gln729. Conversely, the methoxy group on the benzene ring is incorrectly predicted to form specific interactions with the protein, despite being solvent-exposed, though this exposure may facilitate additional binding interactions (blue highlight). Additionally, the pyridine ring, which engages in an arene-arene interaction with Phe729, is partially emphasized by the interpretability model. Notably, although only five atoms of the ligand are among the top 20% atoms, most correspond to distinct protein-ligand interaction sites confirmed by the X-ray crystal structure.
In complex 2WTV (MLN8054 bound to human Aurora kinase A), the core scaffold of MLN8054 comprises a fused polycyclic structure connected to a difluorophenyl group and an anilino moiety, with chlorine and carboxylate substituents. The interpretability of our model effectively highlights key interaction patterns relevant to protein-ligand binding. For instance, within the aniline group, the amino group is correctly identified as forming a hydrogen bond with the backbone of Ala213 (acting as a donor), and the adjacent benzene ring is partially highlighted due to its arene-hydrogen interaction with Gly216. The carboxylate anion is also accurately identified, interacting with Tyr212 via a solvent-mediated contact. Furthermore, the nitrogen-containing aromatic heterocycle is recognized for its potential to act as either a hydrogen bond donor or acceptor, with one specific nitrogen atom explicitly highlighted for donating a hydrogen bond to the backbone of Glu211. The fluorine atoms in the difluorophenyl group and the chlorine substituent are also marked, likely due to their solvent exposure and potential role in enhancing binding affinity. Despite MLN8054 (in 2WTV) being considerably larger and more complex than pyrazoloquinoline (in 3UI7), the model demonstrates strong interpretability by accurately reflecting most of the key interactions confirmed by the crystal structure.
In complex 2QE4 (benzopyran agonist complexed with the human estrogen receptor), the ligand belongs to the class of organic compounds known as 6-hydroxyflavonoids—a subclass of flavonoids characterized by a hydroxyl group at the C-6 position of the flavonoid backbone. The ligand also features several distinctive substituents, including a 1-benzopyran ring, a 1-hydroxy-2-unsubstituted benzenoid moiety, and a 4’-hydroxyflavonoid fragment. Although only a small proportion of the ligand’s atoms are directly involved in protein-ligand interactions, the model accurately identifies these key interaction sites. Notably, the two hydroxyl groups on the aromatic rings are correctly highlighted as forming hydrogen bonds with the receptor, acting as donors to the side chains of Glu353 and His524, respectively. Additionally, two carbon atoms on the five-membered ring are marked as solvent-exposed, suggesting a potential role in mediating further binding interactions.
Regarding the more challenging task of interpreting protein sequences, the results are less specific compared to those for ligands. This limitation arises from the model’s design, where attention weights are computed and visualizations are generated based on protein fragments composed of multiple residues. While this design choice improves computational efficiency, it inevitably reduces interpretability at the residue level, as it does not explicitly resolve which specific residues within each fragment are responsible for binding. Nonetheless, several key amino acid residues involved in binding are correctly highlighted as shown in Figure 4b. For example, in complex 3UI7, residues Gln726 and Phe729, both forming specific interactions with the ligand, are accurately highlighted, along with additional pocket residues such as Tyr524, His525, Leu675, Ser677, and Val678. In complex 2WTV, our model correctly highlights Glu211 and Ala213, which participate directly in ligand binding, as well as several additional binding-site residues: Leu139, Gly140, Lys141, Gly142, Ala160, Leu194, Leu210, and Tyr212. In complex 2QE4, Glu353 and His524 are successfully highlighted as key hydrogen bond partners of the ligand, in addition to pocket-forming residues Thr347, Leu349, Ala350, Met421, Ile424, Gly521, and Leu525. Statistically, across the three complexes, nearly half of the binding-site residues are correctly identified by the model. These results suggest that fragment-level attention provides a promising foundation for interpretability. Incorporating residue-level attention mechanisms could further enhance the resolution and accuracy of protein-ligand interaction explanations. Moreover, if predictions are not constrained to be interaction-free, integrating 3D structural and local spatial information from the binding pocket may further improve interpretability.
Knowledge-level interpretability with knowledge graph
Ablation studies have confirmed that incorporating external knowledge related to proteins and ligands significantly improves affinity prediction. Building on this, we further investigate whether such external knowledge can also enhance the interpretability of protein-ligand binding. To this end, we identify the entities most closely associated with the target protein and ligand and attempt to derive plausible explanations for their interaction based on these entities. Specifically, we utilize the learned embeddings of entities and relations from the KG and continue applying the scoring function to evaluate all triplets involving the target protein or ligand. A lower score indicates greater semantic proximity between the protein or ligand and the corresponding GO or LP entity. As shown in Figure 5, the top five closest entities are presented. These entities are considered the most relevant and are used, along with their definitions, to interpret the protein-ligand binding mechanism. To maintain consistency with prior experiments, we use the same case studies: PDB IDs 3UI7, 2WTV, and 2QE4.
For the PDB structure 3UI7 (pyrazoloquinoline bound to human PDE10A), the top GO entities most closely associated with the protein include cAMP binding, cytosol, 3’,5’-cyclic-GMP phosphodiesterase activity, metal ion binding, and cGMP binding. Definitions for these entities are provided in Figure 5a. Collectively, they suggest that PDE10A functions by catalyzing the hydrolysis of cAMP and cGMP in the cytosol, playing a role in intracellular signal transduction. This enzymatic activity is metal ion-dependent, specifically requiring coordination with $\text{Zn}^{2+}$ and $\text{Mg}^{2+}$ ions. The binding of the ligand pyrazoloquinoline to PDE10A is closely aligned with these biological functions. The ligand binds at the enzyme’s catalytic domain, acting as a competitive inhibitor that blocks access to the natural substrates cAMP and cGMP, thereby preventing their hydrolysis. According to structural data from the PDB, $\text{Zn}^{2+}$ and $\text{Mg}^{2+}$ coordination sites in the catalytic pocket also participate in the protein-ligand interaction. On the ligand side, the top associated entities for pyrazoloquinoline are two aromatic heterocycles, three rotatable bonds, positively charged nitrogen, four hydrogen bond acceptors, and one aromatic carbocycle, with definitions provided in Figure 5b. These descriptors align well with the interpretability analysis shown in Figure 4a. Specifically, the presence of aromatic heterocycles suggests potential for $\pi$-$\pi$ stacking or other aromatic interactions with residues such as Phe, along with hydrogen bonding and structural rigidity, factors known to positively influence binding affinity. Additionally, the ligand’s rotatable bonds may contribute to conformational flexibility, enhancing its ability to adapt to the shape of the protein’s binding pocket. Together, these findings help explain the strong binding affinity observed in complex 3UI7.
For the PDB structure 2WTV (MLN8054 bound to human Aurora kinase A), the top GO entities most closely associated with the protein include protein kinase binding, cytosol, protein serine kinase activity, ATP binding, and protein phosphorylation. These annotations indicate that Aurora kinase A plays a crucial role in spindle assembly and chromosome segregation during mitosis, primarily through its ATP-binding and serine/threonine kinase activities within the cytoplasm. The interaction between the ligand MLN8054 and Aurora kinase A is highly consistent with these functional annotations. MLN8054 is a selective inhibitor of Aurora kinase A that mimics ATP and competitively binds to its ATP-binding site. This binding inhibits the kinase’s activity, suppressing phosphorylation of downstream substrate proteins and thereby disrupting cell division. On the ligand side, the LP entities most closely associated with MLN8054 include acidic group, three aromatic carbocycles, four rotatable bonds, nine heteroatoms, and one aromatic heterocycle. As shown in Figure 4a, the presence of heteroatoms facilitates hydrogen bonding and potential metal coordination, contributing to specific and high-affinity interactions with the protein. Aromatic heterocycles, commonly found in high-affinity drugs, are especially effective in optimizing molecular compatibility with binding pockets. Moreover, a greater number of rotatable bonds enhances the ligand’s conformational flexibility, allowing it to better fit the shape of the binding site. Together, these molecular features help explain the strong binding affinity observed in the 2WTV complex.
For the PDB structure 2QE4 (benzopyran agonist complexed with the human estrogen receptor), the top GO entities most closely associated with the protein are nuclear receptor activity, steroid binding, positive regulation of transcription by RNA polymerase II, estrogen response element binding, and nucleus. Notably, steroid binding highlights the role of the estrogen receptor as a steroid hormone receptor that can be activated by benzopyran agonists exhibiting steroid-like structural features. The remaining GO annotations indicate that, upon activation, the estrogen receptor translocates into the nucleus, binds to specific DNA sequences known as estrogen response elements, and regulates the transcription of target genes involved in processes such as cell proliferation and differentiation. On the ligand side, the most relevant LP entities for the benzopyran agonist in 2QE4 include two hydrogen bond donors, one aliphatic heterocycle, three rotatable bonds, two aromatic carbocycles, and hydrophobe. Among these, the feature of two hydrogen bond donors is consistent with the interaction profile shown in Figure 4a. Hydrophobicity is also a critical factor in protein-ligand binding, as it promotes favorable interactions between nonpolar groups, contributing to the stabilization of the binding site and enhancing overall binding affinity. The implications of the remaining features have been discussed in the earlier examples and similarly contribute to binding specificity and adaptability.
The above examples demonstrate that the KEPLA model can effectively retrieve knowledge entities associated with both the protein and the ligand through KG, thereby leveraging biologically and chemically grounded textual knowledge to explain their binding interactions. Compared to purely data-driven black-box predictions, this knowledge-based interpretability can provide a more tangible and trustworthy foundation for rational drug discovery.
Discussion
In this work, we present KEPLA, an end-to-end knowledge-enhanced deep learning framework for protein-ligand binding affinity (PLA) prediction, designed to address the performance bottlenecks and limited interpretability of existing methods. Specifically, we construct a biochemical KG for proteins and ligands, respectively. The encoded representations of proteins and ligands are integrated into the KG embedding objective to capture the rich semantic information within KG. Subsequently, the model employs cross attention mechanisms to perform fine-grained aggregation of the knowledge-enhanced representations, which are then fed into a decoder to predict binding affinity. We conduct extensive experiments to evaluate the effectiveness of KEPLA in addressing the performance and interpretability challenges.
To assess the predictive performance of our model, we compare KEPLA against several state-of-the-art baselines (spanning both interaction-free and interaction-based methods) on two widely used public datasets. Our experimental setup includes both a random split to assess in-domain performance and a cluster-based split to assess cross-domain generalization. Compared to state-of-the-art baselines, KEPLA consistently achieves superior PLA prediction accuracy under both settings. Notably, despite being an interaction-free model, KEPLA outperforms many interaction-based methods in the in-domain setting and maintains robust performance in the more challenging cross-domain scenario. Ablation study further confirms the critical role of KG in enhancing model performance. These findings suggest that incorporating external domain knowledge not only improves prediction accuracy but also mitigates the hidden data bias inherent in purely data-driven approaches.
To evaluate the interpretability of our model, we conduct case studies that leverage both the cross attention and KG modules to provide explanations from structural and knowledge-based perspectives. By mapping attention weights to protein fragments and ligand atoms, KEPLA can highlight interacting regions such as the protein binding pocket and ligand interacting atoms, offering structural insights into the predicted interactions. Furthermore, by linking proteins and ligands to GO and LP entities in the KG, KEPLA grounds predictions in biochemical context, thereby enhancing interpretability from a knowledge-driven perspective.
This work focuses on interaction-free PLA prediction using 1D protein sequences and 2D molecular graphs as input. Given that high-quality 3D protein structures are available for only a limited subset of known sequences, we deliberately refrain from incorporating 3D structural information in this study. However, recent breakthroughs, such as AlphaFold by DeepMind [48, 49], which has predicted over 2 billion protein structures across a million species, present compelling opportunities for integrating 3D structural information into future chemogenomics-based PLA models. Inspired by principles of pairwise local interaction learning and domain adaptation, we anticipate that extending our approach to include 3D structures could further enhance both predictive performance and model interpretability. Furthermore, our current KG construction is grounded in basic biochemical knowledge. Expanding this to incorporate richer domain-specific and cross-disciplinary knowledge offers a promising direction for future improvement. Finally, the KEPLA framework is highly generalizable and can be readily adapted to other molecular interaction prediction tasks, including drug-drug interaction [50] and protein-protein interaction [51].
Methods
Problem formulation
In PLA prediction, the goal is to estimate the binding strength between a target protein and a ligand compound. We represent each protein sequence as $\mathcal{P}=(a_{1},a_{2},...,a_{K})$, where each $a_{i}$ denotes one of the 23 amino acids. For the ligand compound, most existing deep learning approaches use the SMILES [9], a 1D string format that encodes atoms and bond information in a molecule. While SMILES facilitates the use of classical deep learning approaches, its linear format does not naturally capture molecular structures, potentially leading to the loss of crucial spatial information and reduced predictive accuracy. To address this, our model converts each SMILES string into a corresponding 2D molecular graph. Formally, a ligand is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes (atoms), and $\mathcal{E}$ is the set of edges (chemical bonds) [52]. Given a protein sequence $\mathcal{P}$ and a ligand graph $\mathcal{G}$, the PLA prediction task aims to learn a model $\mathcal{M}$ that maps the joint representation space $\mathcal{P}\times\mathcal{G}$ to a continuous binding affinity value $y\in\mathbb{R}$. The commonly used notations in this paper are summarized in the Supplementary Information.
Overall architecture
As a PLA prediction model, KEPLA consists of four main components: (1) ESM-based protein sequence encoding for learning protein representations; (2) GCN-based molecular graph encoding for learning ligand representations; (3) KG embedding objective, enhancing representations through external biochemical KG; (4) PLA prediction objective, enabling fine-grained modeling of protein-ligand interactions via cross attention. Subsequently, we present a detailed elucidation of these components.
Protein sequence encoding

### v4

@@ Abstract
Accurate prediction of protein-ligand binding affinity is critical for drug discovery. While recent deep learning approaches have demonstrated promising results, they often rely solely on structural features of proteins and ligands, overlooking their valuable biochemical knowledge associated with binding affinity. To address this limitation, we propose KEPLA, a novel deep learning framework that explicitly integrates prior knowledge from Gene Ontology and ligand properties to enhance prediction performance. KEPLA takes protein sequences and ligand molecular graphs as input and optimizes two complementary objectives: (1) aligning global representations with knowledge graph relations to capture domain-specific biochemical insights, and (2) leveraging cross attention between local representations to construct fine-grained joint embeddings for prediction. Experiments on two benchmark datasets across both in-domain and cross-domain scenarios demonstrate that KEPLA consistently outperforms state-of-the-art baselines. Furthermore, interpretability analyses based on knowledge graph relations and cross attention maps provide valuable insights into the underlying predictive mechanisms. The code and datasets of KEPLA are available at the GitHub repository (https://github.com/hanliu95/KEPLA).
I Introduction
The prediction of protein-ligand binding affinity (PLA) is a key task in computational drug discovery [42]. Ligands, including small molecules and biologics, can bind to proteins as agonists or inhibitors. Their interaction strength is measured by binding affinity, which reflects how effectively a ligand can modulate a target protein. Although experimental assays can measure binding affinity, they are labor-intensive and time-consuming. As a result, in silico approaches have attracted increasing attention [34]. These methods enable fast affinity prediction for early-stage drug discovery, facilitating candidate ranking while reducing reliance on costly experimental screening [26].
@@ I Introduction
II Related Work
II-A Protein-Ligand Binding Affinity Prediction
Existing machine learning approaches for PLA prediction can be broadly categorized into two groups based on their core assumptions: interaction-free methods and interaction-based methods.
Interaction-free methods are primarily designed to simplify the modeling process by ignoring the intrinsic 3D structural and interaction information within protein-ligand complexes [4]. For instance, DeepDTA [32] directly employs raw ligand and protein sequences as inputs and utilizes a dual-channel CNN for feature extraction. GraphDTA [31] and MGraphDTA [47] represent ligands as 2D molecular graphs to preserve topological information, while DGraphDTA [13] enhances 1D protein sequence representations through 2D distance maps or protein graphs. Although these models offer computational efficiency by omitting 3D and atomic-level interactions, their generalization ability and interpretability remain limited [3].
In contrast, interaction-based methods explicitly incorporate 3D structural and atomic interaction information as prior knowledge to improve model robustness and interpretability. Early approaches, such as OnionNet [49], typically employ atom-atom pair features for model training, while mathematical tools like topological data analysis have also been introduced for molecular characterization [22, 46]. More recent techniques based on GNNs or 3D-CNNs simplify input representations by using fundamental atomic features such as element types and pairwise distances. For example, Pafnucy [37] adopts 3D-CNNs to extract spatial features from voxelized 3D grids, though their efficiency is hindered by sparse voxel occupancy. Recently, interaction graph neural network (IGNN) models, including SIGN [19] and GIANT [20], have attracted attention for their relational reasoning capabilities, learning intra- and inter-molecular interactions in a staged or type-specific manner. However, these models remain computationally expensive and heavily reliant on accurate structural data, thus limiting their performance in affinity prediction tasks involving novel targets or incomplete interaction information.
II-B Knowledge Graph
KG is a structured semantic network that reveals relations between entities, used to enhance machine cognition and intelligent applications [43]. In recent years, KGs have been gradually introduced into AI4S fields such as protein and molecular research as an effective means of integrating multi-source biological information. By organizing heterogeneous entities and relations—such as proteins, small molecule targets, biological pathways, and experimental data—into structured graphs, KGs provide rich background knowledge and semantic constraints for AI4S tasks [38].
In the field of PLA prediction, existing methods have attempted to jointly model prior knowledge (such as protein families, functional sites, molecular substructures, and pathway relations) with sequence and structural features [27]. By leveraging KG embeddings or GNNs, these methods incorporate global biological context into molecular representations, thereby improving prediction performance compared to relying solely on sequences or structures. Such approaches typically combine KG representations with molecular fingerprints, protein sequence embeddings, or 3D conformational features through concatenation or attention-based fusion, finally obtaining holistic representations that integrate both structural and biological information.
However, existing methods still face limitations: the integration of symbolic knowledge from KGs with physicochemical-based structural features remains nascent, and their robustness is inadequate for scenarios with scarce samples or novel targets. To address these issues, this work proposes an efficient KG-enhanced framework that tightly integrates semantic knowledge with structural information to improve PLA prediction performance.
III Methods
@@
Figure 1: Overview of the KEPLA framework. (a) Existing PLA prediction strategy. (b) KEPLA strategy, which incorporates a biochemical KG to enhance protein and ligand encoding. (c) KEPLA framework with two encoders: an ESM-based encoder for protein sequences and a GCN-based encoder for ligand SMILES, producing local and global representations. Global representations are used for KG embedding, while local representations are fed into the cross attention module for PLA prediction. (d) Cross attention module, where an interaction map derived from protein–ligand local representations enables fine-grained feature fusion, followed by an MLP decoder for affinity prediction. (e) KG embedding module, which projects global representations into a semantic space and evaluates triplet plausibility via a scoring function.
III-A Problem Formulation and Architecture
In PLA prediction, the goal is to estimate the binding strength between a target protein and a ligand compound. Each protein sequence is represented as $\mathcal{P}=(a_{1},a_{2},...,a_{K})$, where $a_{i}$ denotes one of the 23 amino acids. For ligand compounds, although most existing deep learning methods encode molecular information using SMILES strings [44], their linear format does not explicitly capture molecular structure and may lead to the loss of important spatial information. To address this, we convert each SMILES string into a 2D molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents atoms and $\mathcal{E}$ denotes chemical bonds. Given a protein sequence $\mathcal{P}$ and a ligand graph $\mathcal{G}$, the PLA prediction task aims to learn a model that maps $\mathcal{P}\times\mathcal{G}$ to a continuous binding affinity value $y\in\mathbb{R}$.
Based on this formulation, the proposed KEPLA framework is illustrated in Figure 1. KEPLA consists of four main components: (1) ESM-based protein sequence encoding for learning protein representations; (2) GCN-based molecular graph encoding for learning ligand representations; (3) KG embedding objective, enhancing representations through external biochemical KG; (4) PLA prediction objective, enabling fine-grained modeling of protein-ligand interactions via cross attention. Subsequently, we present a detailed elucidation of these components.
III-B Protein Sequence Encoding